OpenStudy (anonymous):

Write the solution in lowest terms. -1/3 + 3/4 = A. 5/12 B. 2/12 C. 1/6

OpenStudy (anonymous):

(-1/3) = (-4/12) (3/4) = (9/12) (-4/12)+(9/12) = 5/12 i feel like I'm doing your math hw for you...

OpenStudy (anonymous):

ya your awesome

OpenStudy (anonymous):

i got 5/6--- wat did i do wrong lol

OpenStudy (anonymous):

answer is A. when doing the greatest common denominator, -1/3is multipied by 4 to get -4/12, then doing that to 3/4. 3/4multiplied by 3 to get 9/12. -4/12+9/12= 5/12

OpenStudy (anonymous):

rofl, i added the top <<< 2 much calc- now i cant even add =( yep ur right guys.